1.) A planet has a diameter of 7.18×106 m and the free fall acceleration on this planet is 6.19 m/s2. What is the mass of this planet?

6.66×1017 kg

1.20×1024 kg

3.33×1017 kg

4.78×1024 kg

2.) A 2.48 kg steel ball and a 1.18 kg wooden ball are 4.36 m apart form center to center. What is the gravitational force that they exert on each other?

1.81×10-11 N

0.671 N

8.70×10-12 N

1.03×10-11 N
Answer #1

q2 =6.67*10^-11 * 1.18 *2.48 / ( 4.36^2 ) =1.02680162e-11 N last option

q1

accleration g = GM/r^2

6.19 = 6.67*10^-11 * x / ( 0.5 *7.18*10^6 ) ^2

x = M = 1.2 *10^24 kg